1. Precision Tool is analyzing two machines to determine which one it should purchase. The company requires a 15 percent rate of return and uses straight-line depreciation to a zero book value over the life of its equipment. Machine A has an initial cost of $892,000, annual maintenance cost of $28,200, and a 4-year life with a market salvage value of $50,000. Machine B costs $1,118,000 initially, has annual maintenance costs of $19,500, and a 5-year life with a market salvage value of $65,000. Both machines will be depreciated straight-line to zero. Whichever machine is purchased will be replaced at the end of its useful life. Assuming a 35% tax rate, Precision Tool should purchase Machine _____ because it lowers the firm's annual cost by approximately _______ as compared to the other machine. Please explain and show your work

Please Not in excel work.
Answer #1

We have following information -

Required rate of rate of return = 0.15

Tax rate = 0.35

Both machine A & B fully depreciated in their useful life

Machine -A Machine-B
Useful life 4 years 5 years
Initial Cost $892,000 $1,118,000
Annual Depreciation (Initial cost/Useful life) $223,000 $223,600
Depreciation Tax shield (Depreciation*Tax rate) $78,050 $78,260
Annual Maintenance cost $28,200 $19,500
Salvage Value $50,000 $65,000

We can calculate Equivalent Annual Cost of Machine in following manner.

NPVofmachine Equivalent Annual Cost(EAC) PVIAF(r, n)

Where,

r = required rate of return

n= useful life

PVIAF = Present value interest annuity factor (cumulative)

Lets calculate Net Present value (NPV) of Machine -A

NPVofAInitialCost Annual Maintenance PVIAF(0.15, 4) DepreciationTarshield PVIAF (0.15,4)- Salvaqevalue PVIF (0.15,4)

PVIAF(0.15,4) = 2.8550

PVIF (0.15,4) = 0.57175

Thus,

NPV of A = 892,000 + 28,200*2.8550 - 78,050*2.8550 - 50,000*0.57175

= 892,000+80,511-222,832.75-28,587.5

= $ 721,090.75

Equivalent Annual cost of A = $ 721,090.75/2.8550

= 252,571.19

NPV of B Initial Cost Annual Maintenance * PVI AF(0.15.5)- DepreciationTarshield * PVIAF (0.15, 5) - Salvagevalue* PVIF(0.15, 5)

PVIAF(0.15,5) = 3.3522

PVIF (0.15,5) = 0.49718

Thus,

NPV of B = 1,118,000 + 19,500*3.3522 - 78,260*3.3522 - 65,000*0.49718

= 1,118,000 + 65,367.9 - 262,343.17 - 32,316.7

= $ 888,708.03

Equivalent Annual Cost of Machine -B = $ 888,708.03/3.3522

= $ 265,111.88

As the Equivalent Annual cost of Machine-B is higher than Machine -A by $ 12,540.69 thus Machine-A is preferable.

Precision Tools should purchase Machine A because it lower's firm's annual cost by approximately $ 12,504.69 as compared to other machine.

NPVofmachine Equivalent Annual Cost(EAC) PVIAF(r, n)

NPVofAInitialCost Annual Maintenance PVIAF(0.15, 4) DepreciationTarshield PVIAF (0.15,4)- Salvaqevalue PVIF (0.15,4)

NPV of B Initial Cost Annual Maintenance * PVI AF(0.15.5)- DepreciationTarshield * PVIAF (0.15, 5) - Salvagevalue* PVIF(0.15, 5)